### v1

@@ 3.2 Large Language Models for visualization
LLM4VIS prompt tuning approaches A novel method for translating natural language queries into visualization is proposed by Chen et al. [4]. The authors involve Codex - a pre-trained LLM developed by OpenAI for natural language to code translation  [^9] https://openai.com/blog/openai-codex - configured within a few-shot learning prompt setup. They propose a two-step strategy to recommend a visualization. Firstly, the input is shaped as examples of natural language to SQL pairs along with the user’s query, letting the model generate the most appropriate SQL. These examples serve as reference points, guiding the model in addressing new user queries effectively. Then, the response generated by the model is translated into Vega-Lite specifications using a rule-based approach [5]. Maddigan et al. [31] present a comparable study involving Codex, GPT-3, and ChatGPT, assessing zero-shot prompting to evaluate the quality of these models on generating visualizations as Python code. A notable limitation of these works is the lack of comprehensive discussion regarding results. Instead, they rely primarily on a small number of qualitative evaluation experiments, merely providing an overview of potential implementations. Li et al. [22] introduce Sheetcopilot, an LLM-based system that uses natural language queries to drive user interaction within spreadsheet software, such as table filtering or chart generation. Sheetcopilot involves in-context learning prompting technique to interact with all spreadsheet functions, employing GPT as model. It utilizes the user’s query alongside a prompt encompassing the model task, a comprehensive list of software functions (i.e., atomic actions), and an output format. With these inputs, Sheetcopilot translates the user request into a systematic plan, outlining step-by-step actions to fulfill the user’s query effectively. A similar approach is Datacopilot introduced by Lin et al.  [24]. Datacopilot is a method to generate automated interfaces and autonomous workflows for human-table interaction, solving different data-related activities, such as acquisition, processing, forecasting, table manipulation, and visualization. It involves GPT in in-context learning to generate and deploy a data processing workflow using small input guidance. However, both Sheetcopilot and Datacopilot, while capable of generating visualization, are not optimized for this task. Their primary focus lies in generating instructions or commands to fulfill user requests within software environments like spreadsheets or toolsets rather than optimizing for insightful visualization.
@@
Overall, while existing prompt tuning approaches offer cost-effective deployment at inference time, they come with limitations. Relying solely on prompt engineering results in a lack of control over the model’s behavior in task execution. Additionally, the majority of these methods depend on proprietary models like GPT. Any updates made to these models may introduce new behaviors, potentially undermining the efficacy of the proposed methodologies.
LLM4VIS fine tuning approaches. A distinct approach to aligning a Large Language Model (LLM) with the visualization task involves fine-tuning the model. Zha et al.  [56] introduce TableGPT, a fine-tuned model designed for querying data tables solely through natural language queries. To mitigate input token length constraints, the authors employ an encoder to preprocess the table, extracting its vector representation. Subsequently, the embedded representation alongside the user query is fed into the LLM, which generates a command set and textual reply. Unlike the previous approach, which focuses on table interactions workflow, Tian et al. [50] propose ChartGPT, an LLM strictly tailored for visualization recommendation. The authors fine-tuned the open-source model FLAN-T5-XL [8], by splitting the data visualization generation problem into two broad tasks (i.e., data transformation and visualization recommendation), each consisting of three sub-tasks. ChartGPT focuses solely on the visualization recommendation task, offering a novel LLM adaptation strategy. However, it lacks a comprehensive visualization narrative. While it can suggest optimal visualizations based on user queries, it fails to explain the reasoning behind the recommendations.
3.3 Generating visualization narratives
@@
Once a model produces a visualization or provides a response to a user query, end users face considerable difficulty in determining the accuracy of the provided output. Compared to textual responses, verifying visualizations poses a more difficult challenge. While tasks such as text summarization, content generation, or code generation allow for relatively straightforward output verification by end-users, the same cannot be said for visualizations because non-expert users may inadvertently consider misleading representations as accurate. Consequently, the production of a ”visualization narrative" (a description of why a recommended visualization fits a user’s intentions implied by the natural language query) emerges as crucial. The task of generating a visualization narrative delivers a textual elucidation of the underlying process behind the visualization generation process performed by the model. This not only enhances transparency but also reduces the potential for misinterpretation or erroneous acceptance of visualized data.
@@
A first approach in this direction has been proposed by Sultanum et al.  [46]. The authors propose Datatales, a method to generate a visualization narrative when given a visualization and an annotation as input. Once the system receives the input, it is passed to GPT-3, which provides a textual narrative based on a simple prompt. Another similar approach is proposed by Ko et al. [19]. The authors present a VL2NL framework to generate a visualization narrative given a VegaLite specification. It accepts as input solely the specification and produces a visualization description and a data description as output. Additionally, it provides a second output, representing a possible query that a generic user could use to generate the visualization. Note that both Datatales and VL2NL perform a different task from NL2VIS, since the visualization is an input, rather than an output.
@@
Unlike all the methods in the literature, we propose V-RECS, a model capable of not only generating a consistent visualization, as proposed by Tian et al. [50], but also explaining the generated output by integrating visualization narrative capabilities, as suggested by Ko et al. [19]. V-RECS aims to combine the strengths of the existing strategies to deliver a system capable of generating and explaining a visualization in a fully-flagged automatic way. Furthermore, V-RECS is designed to be independent of additional software or tools to produce the expected output, as seen in [22, 56, 4]. Instead, we perform all the operations during the reasoning process and provide a verification involving the visualization narrative. V-RECS provides a methodology to adapt smaller models to perform as the larger ones, as proved in the evaluation. Remarkably, V-RECS leverages a ”small" and open-source LLM model, Llama-2-7B, thus mitigating cost, privacy, and controllability issues mentioned in Section 1.
4 Problem formulation
@@
The previous is a general formulation that also applies to NL2VIS systems. LLM4VIS systems introduce a series of intermediate tasks, illustrated in Figure 1. First, it is crucial to note that the model does not directly receive the original data table $T$ as input, but rather $\mathfrak{T}$, which results from a table transformation operation. Given the LLM’s limited input size, depending on the number of input tokens a model can accept, it becomes essential to pre-process the original table into a format suitable for model input. To face this problem, for example, Zha et al. [56] transform the table using an encoder; differently, Tian et al.  [50] propose to use only the table description with some rows to provide the model a better context of the data. Following the table transformation, the input is shaped by a prompt function $P(u_{i},\mathfrak{T})$ (see Section 2) that maps the query and table representation to a designed prompt template.
@@
  • visualization recommendation. Based on the output of the previous steps, the model must choose the best visualization elements (axes, marks, presentation, etc.) to map the data and generate the visualization as code (e.g., matplot python code to produce a bar chart), image (e.g., bar chart as pixel representation), or grammar (e.g., a visualization expressed using VegaLite specifications).
In addition to the above tasks, in this paper we propose that the LLM model should also perform three additional tasks (see the green boxes in Figure 1):
  • visualization caption. The model should generate descriptive captions for the produced visualizations. These captions elucidate the visual representation’s content and meaning, helping the user understand the data shown.
  • visualization explanation.Subsequently, the model should explain why specific columns have been selected from the available dataset. This transparency is essential for users to comprehend the decisions behind the model’s choices, enabling them to verify the relevance and accuracy of the generated visualizations.
  • queries suggestion. Finally, the model should offer suggestions for additional data queries, guiding users on where to delve deeper into the dataset for further insights.
@@
$\centering f_{i}(Q_{i},\mathfrak{T})\rightarrow argmax_{V_{i}}P(V_{i},N_{i}|\mathfrak{T},Q_{i})\quad where\quad V_{i}:\{v_{i,j}|0<j\leq n\}\@add@centering$ (2)
where $N_{i}$ denotes the rich narrative accompanying the generated visualization.
5 V-RECS Model
@@
In this section, we describe the methodology to implement V-RECS. The proposed methodology is based on fine-tuning a small model in a supervised manner, using an existing dataset enriched with visualization narratives generated by a more powerful model. The methodology is depicted in Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions. (the teaser). First (box A, B of Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions.), GPT-4, acting as a Teacher, is instructed to generate visualization narratives leveraging the Chain-of-Toughts technique proposed in [19]. The Teacher receives as input triples from an existing dataset, Ncnet [28], composed by a dataset D, a user’s query Q, and a visualization V matching the user’s needs expressed by Q. Based on the CoT approach, we have defined three tasks, each with a different goal. The first task, T1, aims to explain why a visualization V is a correct answer given the user’s query Q, while the T2 goal involves describing the visualization. Differently, the goal of T3 is to suggest potential areas of interest for further exploration to the user. Next, the teacher-generated responses for the three tasks are combined using a custom prompt template for the subsequent fine-tuning phase (boxes C and D in Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions.).
Then, the responses produced by the larger model are used as ”expertise" to teach a smaller LLM model (the Student LLama-2-7B) to behave similarly (box E). At inference time (right side of Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions.), the fine-tuned model, named V-RECS (Visualization Recommender with Explanations, Captions, and Suggestions), receives as input a (D,Q) pair and recommends a visualization V along with a visualization narrative composed by an explanation (E), a caption (C), and suggestions for further explorations (S). To instruct the models we use the zero-shot prompting strategy (see Section 2) for the Teacher and fully-supervised training for the Student. At inference time instead, V-RECS works in zero-shot. In what follows we explain the methodology in more detail.
5.1 Model training methodology
As described in Section 4, the goal of V-RECS is to recommend visualizations, along with rich narratives to explain and extend recommended visualizations. Furthermore, we also aim to build our solution by leveraging a small and open-source LLM model, the advantages of which we have already described in Sections 1 and 3. Smaller models specialized in a specific task, with comparable performance to larger ones, represent a terrific opportunity to reach the same results at a minor cost, while leveraging an open source software. In the literature, approaches aimed at enhancing the capabilities of LLMs can be broadly categorized into those leveraging the model in its existing form, primarily through prompt tuning, such as Chain-of-Thoughts (CoT) [55], and more advanced techniques involving updating the model weights, like the fine-tuning (see Section 3.2). CoT is one of the most popular methodologies among the former approaches. It pushes the model to emulate human-like cognitive processes by telling it to break down complex problems into more manageable sub-problems, enhancing response accuracy. CoT can be classified as zero-shot [20] and few-shot [55]. The few-shot configuration necessitates a set of question-answer pair examples in addition to the new one. In this case, the prompt typically comprises multiple question-reasoning steps-answer triplets, guiding the model through problem-solving. Conversely, the zero-shot configuration requires a specific text formulation, ’Let’s think step by step,’ within the prompt, telling the model to split the problem into sequential sub-tasks. For instance, a query like ’Q: how much is 2+2? Let’s think step by step’ drives the model to break the question and solve each part individually, culminating in a final response. Unlike the few-shot variant, zero-shot CoT does not provide examples of the problem-solving process but only relies on the model’s embedded knowledge. CoT has demonstrated substantial advantages, particularly in inferring new tasks with minimal effort, without necessitating structural changes to the underlying model architecture [55]. However, its application to models with fewer than 100 billion parameters has been associated with logical inconsistencies during the reasoning process [55]. To address these discrepancies, fine-tuning smaller models emerges as a viable solution to enhance the proficiency of small models in solving complex tasks, such as data visualization. Nonetheless, this approach presents its limitations. Fine-tuning smaller models on complex tasks, using question-answer pairs alone, devoid of reasoning steps, often yields incorrect responses. To overcome these limitations and augment the capabilities of smaller models in data visualization tasks, we propose leveraging a novel approach known as Fine-tuning based on CoT [19]. This method combines both CoT and fine-tuning by enriching question-answer pairs without reasoning steps with the sub-tasks needed to solve the problem posed, improving the reliability of a smaller model in producing correct outputs [1]. The process involves exploiting the embedded knowledge of a larger ’Teacher’ model, such as GPT-4, in a zero-shot CoT configuration, to provide responses enriched with reasoning steps. Subsequently, a smaller ’Student’ model is fine-tuned based on these responses, effectively leveraging the Teacher’s knowledge. For example, for the question: ’A juggler can juggle 16 balls. Half of the balls are golf balls, and half of the golf balls are blue. How many blue golf balls are there?’, the Teacher is tasked with a zero-shot CoT, providing logical steps to produce the expected solution such as: ’Step 1: There are 16 balls in total. Step 2: Half of them are golf balls. Step 3: Therefore, there are eight golf balls. Step 5: etc.’. Then, the response generated and the original question are used to fine-tune the Student model, thereby leveraging the Teacher’s knowledge to enhance the final performance of the generated model.

### v2

@@ III-A LLM4VIS prompt tuning approaches
A novel method for translating natural language queries into visualization is proposed by Chen et al. [31]. Their approach involves a two-step strategy: first, translating the natural language query into SQL, and second, mapping the SQL code to Vega-Lite specifications using a rule-based method [32]. The authors leverage Codex - a pre-trained LLM developed by OpenAI for natural language to code translation [^7] https://openai.com/blog/openai-codex - in a few-shot prompting setup. Codex is provided with examples of natural language to SQL (NL-SQL) pairs, along with the new user query. Based on this input, the model predicts the corresponding SQL, which is then converted into a visualization. Maddigan et al. [33] present a comparable study based on Codex, GPT-3, and ChatGPT to evaluate the quality of these models in generating visualizations as Python code in a zero-shot prompting setup. These papers aim to provide an initial overview of the LLM opportunities in data visualization. However, they lack a thorough quantitative and qualitative evaluation of the proposed methodology. Li et al. [34] introduce Sheetcopilot, an LLM-based framework designed to enhance user interaction with spreadsheet tools like Excel. This framework enables users to make requests in natural language, which are then translated into spreadsheet commands using GPT-4 to achieve the desired outcome. While Sheetcopilot can handle data visualization tasks through chart generation commands, it lacks flexibility and customization in creating visualizations, as it is constrained by the predefined chart structures within the software. A similar approach is Datacopilot introduced by Zhang et al. [35]. Datacopilot generates automated interfaces and autonomous workflows for human-table interaction, addressing various data-related activities such as acquisition, processing, forecasting, table manipulation, and visualization. It leverages GPT to create and implement a data processing workflow with minimal input guidance. While both Sheetcopilot and Datacopilot can create visualizations as an additional feature, they are not optimized for this task. Their primary focus is on deploying interaction workflows and generating instructions or commands to fulfill user requests within software environments like spreadsheets, rather than optimizing for insightful visualizations.
@@
Overall, while existing prompt tuning approaches offer cost-effective deployment at inference time, they come with limitations. Relying solely on prompt engineering results in a lack of control over the model’s behavior in task execution. Additionally, the majority of these methods depend on proprietary models like GPT. Any updates made to these models may introduce new behaviors, potentially undermining the efficacy of the proposed methodologies. An additional concern is privacy. Relying on proprietary models makes it challenging to ensure data protection due to the inability to deploy these models privately.
III-B LLM4VIS fine tuning approaches.
A distinct approach to enhancing a Large Language Model (LLM) for visualization tasks involves fine-tuning the model. Zha et al. [36] introduce TableGPT, an LLM designed to understand and operate on data tables using natural language input. Unlike other approaches, TableGPT relies on a fine-tuned model to predict a set of data analytics commands, such as question answering, data manipulation, visualization, and report generation, based on users’ input. Moreover, Zha et al. propose a novel method to embed large data tables within the limited input length of LLMs. They employ an encoder to preprocess the table, extracting its vector representation. This embedded representation, along with the user query, is then fed into the LLM to predict the optimal command sequence.
While this approach can generate visualizations, it primarily focuses on producing tailored command sequences for performing tasks on data tables. In contrast, visualization recommender systems aim to understand the relationships between queries, data, and optimal visualization configurations, adhering to best practices in data visualization. In line with this objective, Tian et al. [37] propose ChartGPT, a LLM-based model strictly tailored for visualization recommendation. The authors fine-tuned the open-source model FLAN-T5-XL [38], by splitting the data visualization generation problem into two broad tasks (i.e., data transformation and visualization recommendation).
III-C LLM methods for generating visualization narratives
@@
Once a model produces a visualization or provides a response to a user query, end users face considerable difficulty in determining the accuracy of the provided output. Compared to textual responses, verifying visualizations poses a more difficult challenge. While tasks such as text summarization, content generation, or code generation allow for relatively straightforward output verification by end-users, the same cannot be said for visualizations because non-expert users may inadvertently consider misleading representations as accurate. Consequently, the production of a “visualization narrative” (a description of why a recommended visualization fits a user’s intentions implied by the natural language query) emerges as crucial. The task of generating a visualization narrative delivers a textual elucidation of the underlying process behind the visualization generation process performed by the model. This not only enhances transparency but also reduces the potential for misinterpretation or erroneous acceptance of visualized data.
@@
A first approach in this direction has been proposed by Sultanum et al.  [39]. The authors propose Datatales, a method to generate a visualization narrative when given a visualization and an annotation as input. Once the system receives the input, it is passed to GPT-3, which provides a textual narrative based on a simple prompt. Another similar method is proposed by Ko et al. [40]. The authors present a VL2NL framework to generate a visualization narrative given a VegaLite specification. It accepts as input solely the specification and produces a visualization description and a data description as output. Additionally, it provides a second output, representing a possible query that a generic user could use to generate the visualization.
Note that both Datatales and VL2NL are tasked to generate a visualization narrative when provided with a chart or visualization specification. Therefore, they do not belong to the category of visual recommenders.
III-D Our contribution
TABLE I: Comparison of V-RECS and other LLM approaches in the literature. Note that ncNet was included for completeness, as it is based on a Transformer rather than a LLM.
Paper Training strategy Open Weights Model Model Size Visualization Narrative Visualization recommendation Output format
Nl2Interface [31] Prompting $\times$ Codex 12B $\times$ code
Chat2Vis [33] Prompting $\times$ GPT-3 175B $\times$ code
ChartGPT [37] Fine-tuning Flan-t5-xl 3B $\times$ grammar
ncNet [12] Training Transformer n.a. $\times$ grammar
Data-copilot [35] Prompting $\times$ GPT-4 1T E code
TableGPT [36] Fine-tuning Phoneix 7B C commands
SheetCopilot [34] Prompting $\times$ GPT-4 1T $\times$ commands
VL2NL [40] Prompting $\times$ GPT-4 1T ECS $\times$ n.a.
Datatales [39] Prompting $\times$ GPT-3 175B EC $\times$ n.a.
V-RECS Fine-tuning Llama 7B ECS grammar
Table I summarizes the main features of the previously surveyed LLM-based systems. For completeness, we also include NcNet since, although not based on an LLM, it leverages a Transformer architecture. Also note that other traditional NL2Vis approaches are not included in the comparison since, unlike LLMs, they cannot be adapted to generate visualization narratives; rather, they should have been designed with these features from scratch.
To compare the systems with each other and with the solution proposed in this paper, we consider seven dimensions:
  1. 1.
    Training strategy: specifies whether the model uses training from scratch, fine-tuning (pre-trained model aligned to specific tasks), or prompting (tuned at inference time with a prompt).
  2. 2.
    Model, identifies the LLM model used in the proposed strategy (e.g., GPT, Llama).
  3. 3.
    Open Wieghts : Indicates whether the model weights are publicly available (✓) or proprietary ($\times$). Public weights enable private deployment and allow for new fine-tuning strategies, providing more control over the model. In contrast, proprietary models require API interaction and data management on external servers, which can be a limitation for sensitive data and restrict users from using the provided methods to enhance/refine the model.
  4. 4.
    Model size: Along with open weights, the model size is crucial for private deployment. Smaller models like Llama 7B can be easily deployed on commercial computers, while larger models require more computational resources, posing a limitation for private deployment.
  5. 5.
    Visualization recommendation : this dimensions tells if the system performs (✓) or does not perform ($\times$) the task.
  6. 6.
    Visualization narrative : this dimension specifies which types of narratives are generated (E, C, S), while $\times$ means that no narratives are produced.
  7. 7.
    Output Format: this column compares visualization methods based on how they generate visual outputs. Some methods provide visualization as code or grammar, allowing for highly customizable and flexible visual representation. On the other hand, tools that generate predefined commands for visualizations offer a more guided -but less flexible- approach, where users utilize pre-packaged commands to create visual outputs. Note that, since VL2NL and Datatales do not generate visualizations, this classification is not applicable (n.a.).
@@ III-D Our contribution
Hereafter, we propose V-RECS, a model capable not only of generating a coherent visualization, as proposed by Tian et al. [37], but also of explaining the generated output by integrating narrative visualization capabilities, as suggested by Ko et al.  [40]. Unlike the few systems that produce visualization narratives in the form of captions, V-RECS generates a very rich narrative, composed of a detailed explanation of the reasoning process that led to the production of the visual recommendation, a caption to describe the recommended chart, and suggestions for further exploration of the data. Furthermore, V-RECS leverages a methodology to adapt smaller models to perform as the larger ones, as demonstrated in Section VI. Remarkably, V-RECS’ fine-tuning strategy exploits a “small” and open-source LLM model, thus mitigating cost, privacy, and controllability issues mentioned in Section I.
IV Problem formulation
@@
The previous is a general formulation that also applies to NL2VIS systems. LLM4VIS systems introduce a series of intermediate tasks, illustrated in Figure 1. First, it is crucial to note that the model does not directly receive the original data table $T$ as input, but rather $\mathfrak{T}$, which results from a table transformation operation. Given the LLM’s limited input size, depending on the number of input tokens a model can accept, it becomes essential to pre-process the original table into a format suitable for model input. To face this problem, for example, Zha et al. [36] transform the table using an encoder; differently, Tian et al.  [37] propose to use only the table description with some rows to provide the model a better context of the data. Following the table transformation, the input is shaped by a prompt function $P(u_{i},\mathfrak{T})$ (see Section II) that maps the query and table representation to a designed prompt template.
@@
  • visualization recommendation. Based on the output of the previous steps, the model must choose the best visualization elements (axes, marks, presentation, etc.) to map the data and generate the visualization as code (e.g., matplotlib python code to produce a bar chart), image (e.g., bar chart as pixel representation), or grammar (e.g., a visualization expressed using VegaLite specifications).
In addition to the above tasks, in this paper we propose that the LLM model generates a visualization narrative VN, which we split into three tasks (see the green boxes in Figure 1, and the right-bottom side of Figure 1 for an example):
  • Explanation (E). First, the model should explain why specific columns have been selected from the available dataset, and which transformation functions (e.g., average) have been applied. This transparency is essential for users to understand the decisions behind model choices, allowing them to verify the accuracy of the generated visualization and its adherence to users’ information needs.
  • Captioning (C). Subsequently, the model should generate descriptive captions for the produced visualizations. These captions elucidate the visual representation’s content and meaning, helping the user understand the data shown.
  • Suggestion (S). Finally, the model should offer suggestions for additional data queries, guiding users on where to delve deeper into the dataset for further insights.
@@
$\centering f_{i}(Q_{i},\mathfrak{T})\rightarrow argmax_{V_{i}}P(V_{i},VN_{i}|\mathfrak{T},Q_{i})\quad where\quad V_{i}:\{v_{i,j}|0<j\leq n\}\@add@centering$ (2)
where $VN_{i}$ denotes the rich narrative accompanying the generated visualization.
V V-RECS Model
@@
In this section, we describe the methodology to implement V-RECS. The proposed methodology is based on fine-tuning a small model in a supervised manner, using an existing dataset enriched with visualization narratives generated by a more powerful model. The methodology is depicted in Figure 2. First (box A of Figure 2, left), GPT-4, acting as a Teacher, is instructed to generate visualization narratives leveraging the Chain-of-Toughts zero-shot prompting technique proposed by Ko et al. [40] (see Section II). The Teacher receives input from an existing dataset composed of triples (D,Q,V) where D is tabular data, Q a user’s query in natural language, and V a visualization specification matching the user’s needs expressed by Q. Based on the CoT approach, the Teacher is prompted with three CoT tasks, each with a different goal. The first task, T1, is to explain why the visualization V is a correct answer, given the user’s query Q, while the goal of T2 involves generating a caption to describe the chart. Finally, the goal of T3 is to suggest potential areas of interest for further data exploration to the user. The Teacher, based on the three prompts, generates the responses for the three tasks, $R_{1}$, $R_{2}$, and $R_{3}$ (box B). Next, the responses are combined, along with the original triple (D,Q,V), to create training instances for the subsequent fine-tuning phase (box C). During fine-tuning (box D), the training instances - which are now composed of sextuples (D,Q,V,E,C,S) - are used to teach a smaller LLM model (the Student LLama-2-7B) to generate visual recommendations along with visualization narratives. At inference time (right-hand side of Figure 2), the fine-tuned model, named V-RECS (Visualization Recommender with Explanations, Captions, and Suggestions), receives as input a (D,Q) pair and recommends a visualization V along with a narrative VN composed by an explanation (E), a caption (C), and suggestions for further explorations (S).
To instruct the models we use the zero-shot prompting strategy for the Teacher and fully supervised training to fine-tune the Student. At inference time instead, V-RECS works in zero-shot. In what follows we explain the methodology in more detail.
Figure 2: During the training phase (left), a large LLM, GPT-4, acting as a Teacher (A), receives as input triples (D,Q,V) extracted from an available dataset, where D is tabular data, Q is a natural language query, and V is the VegaZero specification for a chart matching the user’s query Q. The model is prompted to perform three Chain of Thought (CoT) tasks: T1, to explain the reasoning steps that justify the answer V given the query Q; T2, to generate a caption that describes the chart, and T3, to suggest additional useful queries. The responses of the model are combined to create a visualization narrative (VN) used to augment the initial dataset (C). Next, a small model, Llama-2-7B (the Student), is fine-tuned (D) with the enriched dataset of quadruples (D,Q,V,VN). At inference time (right), the resulting specialized model, named V-RECS, receives as input a pair (D,Q) and generates, along with a visualization recommendation V, a visualization narrative VN made of an explanation E, a caption C, and a suggestion S of additional queries. Adopting a teacher-student metaphor, the esteemed professor GPT-4, leveraging its trillion-parameters knowledge of data visualization and natural language understanding, generates examples of visualization narratives for the young student Llama-2-7B. Llama uses this supplementary material to diligently learn the task of producing similar narratives alongside the recommended visualization, and finally gets its doctorate in Visual Recommendation with Narratives, becoming Dr. V-RECS. During the exercise of its profession, will Dr. V-RECS be able to match Professor GPT-4’s abilities? Certainly, its professional fees are much cheaper.
V-A Model training methodology
As described in Section IV, the goal of V-RECS is to recommend visualizations, along with rich narratives to explain and extend recommended visualizations. Furthermore, we also aim to build our solution by leveraging a small and open-source LLM model, the advantages of which we have already described in Sections I and III. Smaller models specialized in a specific task, with comparable performance to larger ones, represent a terrific opportunity to reach the same results at a minor cost while leveraging an open source model.
To enrich the explainability capabilities of our model and instruct it to generate rich visualization narratives, we leverage Chain-of-Thoughts, a reasoning approach that mimics human problem-solving by breaking down complex tasks into simpler, sequential stages (see Section II). CoT has demonstrated substantial advantages, particularly in inferring new tasks with minimal effort, without necessitating structural changes to the underlying model architecture [21]. However, its application to models with fewer than 100 billion parameters has been associated with logical inconsistencies during the reasoning process [21]. To address these discrepancies, fine-tuning smaller models emerges as a viable solution to enhance the proficiency of small models in solving complex tasks, such as data visualization. Nonetheless, this approach presents its limitations. Fine-tuning smaller models on complex tasks using question-answer pairs alone, devoid of reasoning steps, often produces incorrect responses.
To overcome these limitations and augment the capabilities of smaller models in data visualization tasks, we propose leveraging a novel approach known as Fine-tuning based on CoT [40]. This method combines both CoT and fine-tuning by enriching question-answer pairs without reasoning steps with the sub-tasks needed to solve the problem posed, improving the reliability of a smaller model in producing correct outputs [41]. The process involves exploiting the embedded knowledge of a larger “Teacher” model, such as GPT-4, in a zero-shot CoT configuration, to provide responses enriched with reasoning steps. Subsequently, a smaller “Student” model is fine-tuned based on these responses, effectively leveraging the Teacher’s knowledge.